In a simple random sample of size 98, there were 64 individuals in the category of interest. It is desired to test H0: p = 0.73 versus H1: p < 0.73. Compute the test statistic z. A) -1.72 B) 0.04 C) 6.19 D) 0.65
Answer #1

solution:

= x / n = 64/98=0.653

P0 = 0.73

1 - P0 = 1-0.73=0.27

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.653-0.73 / [(0.73*0.27) / 98]

= -1.72